OpenStudy (anonymous):

an engineer works at a plant out-of-town. a car is sent for him from the plant everyday that arrives at the railway station at the same time as the train he takes. one day the engineer arrived at the station one hour before his usual time and , without waiting for the car ,started walking for work . on his way he met the car and reached the plant 10 minutes before the usual time. how long did the engineer walk before he met the car ?

OpenStudy (anonymous):

in units it should be 6

OpenStudy (anonymous):

you have to ans ... for how much time he had to walk.

OpenStudy (anonymous):

then w8

OpenStudy (anonymous):

nw see fr d usual routine if you plot along the time horizon scale.at t=0 he is at the station and car is there to pick up

OpenStudy (anonymous):

bt when he arrives 1 hour before..then shift the time horizon

OpenStudy (anonymous):

ans is 55 minutes..

OpenStudy (anonymous):

@shreya pls share how do u find that???????//

OpenStudy (anonymous):

well you can start like this- suppose the train and car reach at 10 am but one day the train arrived at 9 am and the engineer started to walk ... you'll get the answer if u go this way... i m stuck ..i just need some help to get the ans. i have done it before but now m not able to get the ans..

OpenStudy (shubhamsrg):

Lets say the train used to arrive at T . it took 10 mins less, thus the car must have picket the man 5 mins before T,symmetrically. Hence we walked for 60-5 mins = 55 mins

OpenStudy (anonymous):

@shubhamsrg can u explain in detail...

OpenStudy (shubhamsrg):

It was a good question, hmm

OpenStudy (anonymous):

hey why did yu delete it ???

OpenStudy (shubhamsrg):

Understand it like this: lets say his train arrives at 12:00 (i.e. T) On a usual day, car will leave plant at 12:00 - x , will arrive at station at 12:00 reaches work at 12:00 + x on this very special day, man starts walking at 11:00 , car leaves plant at same 12:00 - x car meets man at 12:00 - x + y He reaches to work at 12:00 - x + y + y Note that 12:00 - x+2y = 12:00 + x - 10 => x-y= 5 Thus car meets man at 12:00 - x + y = 12:00 - 5 = 11:55 Hence he walked for 55 mins

OpenStudy (shubhamsrg):

nevermind.

OpenStudy (anonymous):

ah..!! brilliant ..!! amazing yar.. thanks :D .

OpenStudy (shubhamsrg):

Its one of those rare occasions when I did a question! :P Thanks anyways. :)

OpenStudy (anonymous):

bt you can also use shifting the time horizon as we do in our transformation of curves

OpenStudy (anonymous):

@manishsatywali I need an easy explanation..i have no idea about what you're saying

OpenStudy (shubhamsrg):

Firstly I was going all crazy on it with too many variables. I even tried graphing but failed. This one some how worked out.

OpenStudy (anonymous):

k i got it

OpenStudy (anonymous):

u will learn transformation of curves............same concept applies here